OpenStudy (anonymous):

some one please tell me what to write this grade could save me from retaking the class 1. [9.01], [9.02] Part 1: Solve each of the quadratic equations below and describe what the solution(s) represent to the graph of each. Show your work to receive full credit. x2 − 25 = 0 and x2 = 2x + 15 Part 2: Using complete sentences, compare and contrast the graphs of y = x2 − 25 and y + x2 = 2x + 15.

OpenStudy (triciaal):

x2 − 25 = 0 difference of 2 squares x^2 - 5^2 a^2 - b^2 = (a + b)(a-b)

OpenStudy (anonymous):

for some reason it wont let me see any thig behind x2 and betwen 25 =0

OpenStudy (triciaal):

x2 = 2x + 15 x^2 -2x -15 = 0

OpenStudy (triciaal):

refresh the page

OpenStudy (anonymous):

got it what abuut part two sorry im lost in this lesson

OpenStudy (triciaal):

it reads x^2 -25 = 0

OpenStudy (triciaal):

look above then continue here

OpenStudy (anonymous):

is it lets me see now

OpenStudy (triciaal):

what are the factors of 15 that add up to -2 ( pay attention to the signs)

OpenStudy (anonymous):

-5 and 3

OpenStudy (triciaal):